the length of a rectangular poster is 5 inches longer than the width if. the permeter of the poster is 122 inches what is the width?
Answer by TimothyLamb(4379)

x = y + 5
122 = 2x + 2y
---
put the system of linear equations into standard form
---
x - y = 5
2x + 2y = 122
---
copy and paste the above standard form linear equations in to this solver:
https://sooeet.com/math/system-of-linear-equations-solver.php
---
answer:
length = 33 in
width = 28 in
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php